OpenStudy (andrewkaiser333):

Joe's test grades in History class are: 92%, 89%, 85%, 89%, and 90. The semester final will count as two tests. Joe needs to get a grade of 90% or higher for the semester to get an A. What is the minimum score Joe can get on the final test and average 90%?

OpenStudy (mathstudent55):

How do you find the average grade of several grades?

OpenStudy (andrewkaiser333):

can you look it up i am stumped i was gone from school for several days due to the death of my mom

OpenStudy (andrewkaiser333):

I tried and i found nothing i am not sure what you look for first

OpenStudy (mathstudent55):

To find the average of several grades, add up all the grades, and divide by the number of grades.

OpenStudy (andrewkaiser333):

ok can you start the equation?

OpenStudy (mathstudent55):

You are given these test grades: 92, 89, 85, 89, 90 He will take a final exam and get one more grade that counts like two tests. Since we don't know what his final exam grade is, we call it x.

OpenStudy (mathstudent55):

That means the sum off all the grades will be: \(92+ 89+ 85+ 89+ 90 +x + x\) Ok so far?

OpenStudy (andrewkaiser333):

ok

OpenStudy (mathstudent55):

To find the average, you divide the sum of the grades by the number of grades. There are 7 grades altogether, so to find the average, we divide the sum by 7. \(\dfrac{92+ 89+ 85+ 89+ 90 +x + x}{7}\) Ok?

OpenStudy (andrewkaiser333):

k

OpenStudy (mathstudent55):

We want the average to be 90 or more, so we now set it up as an inequality: \(\dfrac{92+ 89+ 85+ 89+ 90 +x + x}{7} \ge 90\)

OpenStudy (mathstudent55):

Now we need to solve the inequality for x.

OpenStudy (andrewkaiser333):

k

OpenStudy (mathstudent55):

First, add all the numbers on the numerator of the fraction. Also, what is x + x = ?

OpenStudy (andrewkaiser333):

\[\frac{ 445+2x }{ 7 }\

OpenStudy (andrewkaiser333):

\[\frac{ 445+2x }{ 7 }\]

OpenStudy (andrewkaiser333):

yes?

OpenStudy (mathstudent55):

Great, so now we have this: \(\dfrac{445 +2x}{7} \ge 90\)

OpenStudy (andrewkaiser333):

now x7

OpenStudy (mathstudent55):

Correct. Now multiply both sides by 7 to get rid of the denominator of 7. \(7 \times \dfrac{445 +2x}{7} \ge 7 \times 90\)

OpenStudy (andrewkaiser333):

630

OpenStudy (andrewkaiser333):

445+2x >630

OpenStudy (andrewkaiser333):

-445

OpenStudy (andrewkaiser333):

2x>185

OpenStudy (mathstudent55):

\(\cancel{7} \times \dfrac{445 +2x}{\cancel{7}~1} \ge 630\) \(445 + 2x \ge 630\) Now subtract 445 from both sides.

OpenStudy (andrewkaiser333):

/2

OpenStudy (mathstudent55):

Correct. Now divide both sides by 2.

OpenStudy (andrewkaiser333):

92.5>x

OpenStudy (mathstudent55):

No. Be careful. Don't switch sides.

OpenStudy (andrewkaiser333):

oops

OpenStudy (andrewkaiser333):

how is it done then

OpenStudy (andrewkaiser333):

x>92.5

OpenStudy (mathstudent55):

We had \(2x \ge 185\) Divide both sides by 2: \(\dfrac{2x}{2} \ge \dfrac{185}{2}\) We get: \(x \ge 92.5\)

OpenStudy (andrewkaiser333):

ok

OpenStudy (mathstudent55):

Correct, but remember it's "greater than or equal", not just plain "greater than."

OpenStudy (andrewkaiser333):

thanks i see now

OpenStudy (andrewkaiser333):

yes i know

OpenStudy (andrewkaiser333):

my keys just won't let me click to do it the way you did it

OpenStudy (andrewkaiser333):

it stopped working for some reason

OpenStudy (mathstudent55):

Since the answer is \(x \ge 92.5\) That means as long as he gets at least 92.5% on the semester exam, he will have an 90% average, meaning he gets an A grade.

OpenStudy (mathstudent55):

You're welcome.